OpenStudy (anonymous):

The slope of a line is -, and the y-intercept is 5. What is the equation of the line written in general form? A) x - 5y - 25 = 0 B) x - 5y - 5 = 0 C) x + 5y - 25 = 0 Quick answer please!!

OpenStudy (anonymous):

copy and paste from exam fail

OpenStudy (anonymous):

um what?

OpenStudy (anonymous):

The slope of a line is -, and the y-intercept is 5. What is the equation of the line written in general form?

OpenStudy (quietus):

Please follow @tkhunny's answer from your previous question... Anytime you are given the slope (or a hint about the slope) and the y-intercept, the Slope-Intercept is normally the way to go. A) 10x + 3y - 1 = 0 B) x + 3y + 10 = 0 C) x + 3y - 10 = 0 Move the constant with subtraction. A) 10x + 3y = 1 B) x + 3y = -10 C) x + 3y = 10 Move the x-term with subtraction. A) 3y = -10 x + 1 B) 3y = -x - 10 C) 3y = -x + 10 Isolate the the 'y' and produce the final Slope-Intercept Form. A) y = -(10/3) x + 1/3 B) y = -(1/3)x - 10/3 C) y = -(1/3)x + 10/3 Identify the parts A) y = -(10/3) x + 1/3 <== Slope: -10/3 and y-intercept 1/3 B) y = -(1/3)x - 10/3 <== Slope: -1/3 and y-intercept -10/3 C) y = -(1/3)x + 10/3 <== Slope: -1/3 and y-intercept 10/3 Find the one that fills all the requirements. Be deliberate and systematic. No need to guess or wonder.

OpenStudy (quietus):

Please do not just ask for quick answers. It is important that you understand the process so that you can do the questions on your own.

OpenStudy (anonymous):

not really much chance of anyone answering The slope of a line is -, and the y-intercept is 5. What is the equation of the line written in general form? in any case

OpenStudy (anonymous):

im in summer school just trying to make up credits from freshman year, and I'm a senior now, I'm just worried about getting it done before monday! @quietus

OpenStudy (anonymous):

ok i will say it once more and then disappear this is what you wrote The slope of a line is -, and the y-intercept is 5. What is the equation of the line written in general form? missing something?

OpenStudy (anonymous):

ohhhhh the slope of the line is -1/5 @satellite73

OpenStudy (quietus):

actually the slope isn't very important in this case because its a multiple choice question. As long as you know its negative, it means that x is negative.

OpenStudy (quietus):

and knowing that x is negative, A and B are immediately ruled out.